In my undergraduate Statistics class, the professor called on students to answer questions using a true random number generator (all events were independent). I could never be sure if I was going to be called on next! He would just hit a button and a name would appear on the screen! Yikes! For the remaining problems, assume that there were 25 students in my Statistics course.

What is the probability that a student would get called on to answer BOTH of the first two questions in class?
Answer #1

Probability of a student being called = 1/25.

So, the probability that a student would get called on to answer both of the first two questions = (1/25)*(1/25) = 1/625 = 0.0016